A father is 5 times as old as his son. In 18 years, the father will be twice as old as his son. How is old is the father at present?

Answer by richwmiller(17219)

f=5s,
f+18=2*(s+18)
5s+18=2*(s+18)
5s+18=2s+36
3s=18
s=6
f=30
dad is 30 now